OpenStudy (anonymous):

What is the simplified form for the product x+1/x^2-25 times x+5/x^2+8x+7

OpenStudy (ajprincess):

First factorise x^2-25 and x^2+8x+7 x^2-25=(x+5)(x-5) x^2+8x+7=x^2+7x+x+7 =x(x+7)+1(x+7) =(x+1)(x+7)